OpenStudy (anonymous):

If 12 ft2 of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (anonymous):

??

OpenStudy (ddcamp):

Let one side of the base be L. Area of the base = L^2 Area of each side = (12-L^2)/4 Height = (area of side / width of side) = (12-L^2)/(4L) Volume = B×h = (L^2) × (12-L^2)/(4L) = L×(12-L^2)/4 \[V = L(12-L^2)/4 \\ = -\frac{ 1 }{ 4 } L^3 + 3L\] Now, what's the highest V can be (knowing that L is between 0 and sqrt 12)